If a,b,c be the unit vector such that b is not parallel to c and a×(2b×c)=b then angel that a make with b and c respectively is ?

Grade:12

1 Answers

Aditya Gupta
5 years ago
the equation becomes:
2[(a.c)b – (a.b)c]=b
b[2(a.c) – 1]= (a.b)c
so, a.b=0 and 2(a.c) – 1=0 or cosθ=1/2=cos60
so, angle between and is 90 degree
and angle between  and is 60 degree